4.Create a scatter plot using the data provided below (Note. This includes labeling the correct axis for X and Y). Also, identify what type of correlation was observed (if any). Aggression (X) 5,3,1,6,8Self-Esteem (Y)4,2,3,8,7
Answer #1

Here we see that there is increasing trend which means for every increase in x, there is corresponding increase in y.

Hence there is positive correlation between x and y.

Let us also compute value of r

X Values
∑ = 23
Mean = 4.6
∑(X - Mx)2 = SSx = 29.2

Y Values
∑ = 24
Mean = 4.8
∑(Y - My)2 = SSy = 26.8

X and Y Combined
N = 5
∑(X - Mx)(Y - My) = 22.6

R Calculation
r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy))

r = 22.6 / √((29.2)(26.8)) = 0.8079